1)Write an equation of each line in standard form
y=-3/2x-1/4
2) Write an equation for the line
m=2/5 and the y intercept is 17/5

Answer by rothauserc(4718)

standard form of a line is
:
Ax + By = C
:
where A, B, C are real numbers
:
1) y =-3/2x -1/4
:
multiply both sides of = by 4
:
4y = -6x -1
:
6x + 4y = -1
:
2) y = 2x/5 + 17/5
;
multiply both sides of = by 5
:
5y = 2x + 17
:
2x - 5y = -17
: